The length of a rectangle is 6 meters longer than the width. If the area is 40 square meters, find the rectangle’s dimensions. Round to the nearest tenth of a meter.
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The length of a rectangle is 6 meters longer than the width. If the area is 40 square meters, find the rectangle’s dimensions.
======================
Find a pair of factors of 40 that differ by 6.
-----
1*40 NG
2*20 NG
5*8 NG
etc

Answer by ikleyn(52937)

.

1.  "Trial and error" solution (by guessing) is 4 meters (width) and 10 meters (length) and takes 4 seconds.



2.  Algebra solution is via this equation


        x*(x+6) = 40

        x^2 + 6x - 40 = 0

        (x+10)*(x-4) = 0


    giving the same answer x= 4 meters for the width  and  (x+6) = (4+6) = 10 meters for the length.

Solved.